# Task Facet Learning: A Structured Approach to Prompt Optimization

### v1

@@ 4.3 Results on a Real-world Task: Search Query Intent
We consider the task of inferring if two search queries share identical intent or not. This is an important task arising in search and recommendation pipelines, and is tackled today using LLMs. It is challenging to tell if two queries share the same intent, as it requires domain knowledge (e.g., brands and product categories), depends on geographical biases (e.g., “cricket” and “cricket game” are likely to mean the same in UK, but unlikely in the US), and even on cultural connotation. So, examples are crucial for understanding the task and engineering a prompt that generalizes well.
@@
The prompt obtained using UniPrompt improves over the best manual prompt by 5.77% on the negative (rare) class, by 0.23% on the positive class, and by 1.86% overall on the test set. The learnt prompt captures the following facets of the task: (1) recognizing variations in names and abbreviations, and how they do not necessarily change the context; (2) recognizing the specificity of brands, and how even minor variations do change the context; and (3) recognizing the specificity of terms in queries, and how lack of specific terms can indicate departure of intent.
4.4 Do diverse task facets organized as sections really help?
We now revisit our hypothesis and motivation in Section 2.2. We want to empirically validate if all the diverse task facets that UniPrompt learns indeed contribute to the performance gains that we observe in Table 1. We consider two ablations:
4.5 Ablations: Impact of Clustering, Inclusion of History, and Greedy Update
Here we study effect of our design choices, namely clustering (Section 3.2), inclusion of history of past prompts (Section 3.2), and the greedy update rule (Section 3.2.3). The results are shown in Table 3. We see that clustering as well as edit history components are critical for performance of UniPrompt in all the datasets. We see a major drop of $14.8\%$ in accuracy in the Ethos dataset when clustering is removed, and a $4.3\%$ drop when history component is removed. In all the datasets except GSM8K, we find clustering is more important than history. This can attributed to variability of question types being less in GSM8K than in others: all questions are grade-8 arithmetic questions in GSM8K.
We also find that the greedy update rule proves to be superior or competitive compared to beam search in relatively easier datasets — where even less exploration produces good results, greedy proves to be a more effective update rule. On the other hand, in more complex datasets like MedQA, greedy appears to be a bad strategy. We also see that clustering examples based on feedback (“Fb Clustering”) is a better strategy than clustering based on topics, except for the Ethos dataset.
5 Related Work
Here, we highlight relevant work that are not addressed in the manuscript so far. Deng et al. [6] present a discrete prompt optimization method, RLPrompt, using reinforcement learning, where a policy network learns to generate effective prompts through reward-based training, with an emphasis on enhancing training efficiency through effective reward stabilization techniques. A drawback of such automatic prompt optimization approaches [22, 36, 6, 31] is that the prompts generated tend to be short, often comprising only one or two sentences, which may not fully encapsulate the complexity of the task at hand.
Prior research [28, 27] has highlighted the significance of specific sections within prompts. For instance, a well-designed prompt typically incorporates elements such as a system prompt and a chain-of-thought section. However, existing methods do not specifically target the optimization of individual sections and their respective contents within the prompts. Hsieh et al. [11] investigate the use of greedy and genetic algorithms to edit lengthy prompts. Their method focuses on paraphrasing one line at a time starting from an existing prompt, compared to our goal of learning facets of a task from scratch, thereby compromising generalization accuracy. Another orthogonal line of work that ha been explored recently for prompt optimization is the algorithmic selection of in-context examples [19, 9, 30, 25, 26].
6 Limitations and Future Work
We presented a method inspired by the human prompt engineering process to generate complex prompts from scratch that include different facets of a task. Our algorithm provides significant improvements over baseline prompt generation methods on multiple standard datasets. However, in our evaluation, we considered only classification tasks and only GPT-4 as the expert LLM. Generalizing to generative tasks, as well as investigating the impact of the expert LLM choice are some future directions.
Appendix A Appendix
Ours Ours + KNN Ours + KNN + COT Ours + KNN+COT+Ensemble MedPrompt
MedQA 80.9 81.0 83.9 87.0 80.6
PubMedQA 70.3 72.2 74.7 75.6 71.2
MedMCQA 79.2 81.4 82.6 84.5 79.1
MMLU MG 78.0 94.0 96.0 99.0 98.0
Figure 3: Training curves for MedQA (left) and ARC (middle) datasets when UniPrompt is initialized with (published) state-of-the-art prompts; (right) ablation of facets on Ethos (Section 4.4).
Ethos ARC MedQA GSM8K
UniPrompt -- History 88.0 84.6 55.3 80.8
UniPrompt -- Clustering 77.5 82.0 54.1 81.5
UniPrompt 92.3 86.0 57.1 82.4
UniPrompt $+$ Greedy 93.7 90.5 55.5 82.3
UniPrompt $+$ Fb Clustering 87.2 91.2 58.3 82.5
A.1 Details on estimation of Lipschitz constant $L$
TO calculate the Lipschitz constant for a given LLM and task, we take a human written promp and generate it’s paraphrases using GPT-4. We prompt GPT-4 with the following text: “You are given a sentence, you have to generate 30 paraphrases of the sentence, make sure that the cor content of each paraphrase is same, you can use add, subtract or change words". These paraphrases are then evaluated on the validation set $D_{v}$. For a measure of distance between two prompts, we take the cosine similarity between the embeddings of two prompts. We use text-ada-002 for generating the text embeddings for prompts.
A.2 Details on marginal utility difference calculation for prompts
For a given LLM and task, to study the marginal utility of different facets for a prompt, we first sample a few facets (generated using the fine-tuned model) that individually perform better than the chain-of-thought prompt as a baseline. Given these set of facets, we sample three facets $f_{1},f_{2},f_{3}$ and generate four prompts: $A:f_{1},B:f_{2},A+\Delta:f_{1}+f_{3},B+\Delta:f_{2}+f_{3}$. We define $\Delta_{A}$ as the accuracy difference between prompt $A$ and prompt $A+\Delta:f_{A}+f_{A+\Delta}$. Similarly, $\Delta_{B}=f_{B}+f_{B+\Delta}$
A.3 Prompt Initialization
@@ Introduction:
Assume the role of a science expert and answer the given
question by selecting one of the options A, B, C or D.
         1. Understand and solve science questions by selecting
         the best answer from a given list of options.
         2. Identify the logic behind the choices provided and
         make an informed decision.
         3. Use contextual clues to choose the most accurate
         answer.
         4. Be aware of the differences between science and
         everyday language.
@@ 4. Look for clues: Look for clues in the question and the
A.4 SLM Training Details
To induce the ability of structured prompt generation in a smaller language model, we curate a section-wise dataset of around 12,000 task-prompt pairs. The tasks for training dataset creation were taken from tasksource library [24] that contains around five hundred classification tasks. We extract the task description from tasksource-instruct, which contains tasksource dataset recasted with instructions. For instance, the task description for BIG-bench Entailed Polarity task is, "Given a fact, answer the following question with a yes or a no". The dataset provides diverse tasks and their short description, but not the human-generated prompts for each task. To approximate human-generated prompts, we use GPT-4 as a teacher model.
A.5 Prompt to Llama2-13B for fine-tuning
### Instruction:
You are a prompt engineer, you have to write a structured prompt.
For the given task description, examples and section description,
write the contents of the section that align with section description.
### Task Description:
{data_point[’task_description’]}
### Section Description:
{data_point[’section’]}:{section_descriptions[data_point[’section’]]}
### Response:
{data_point[’prompt’]}
A.5.1 Examples of training dataset
@@ 3. To make the task more manageable, first focus on swapping the
A.6 Prompt for identifying important facets
@@ A.9 Feedback Prompts
You are a teacher and you have to give feedback to your students on their answers.
@@
You are given a question, it’s true answer and answer given by student.
You are also given the explanations written by your students while solving the questions.
@@
which edit can be applied to aid the student in solving the wrong question.
@@ 5. You can add specifics like examples, definitions etc make
Your summary cannot include more than four functions. Make sure that the content is useful,
@@ A.10 Editing Prompt
A.11 Example of prompt evolution using our method
@@
A.12 Comparision of our method with existing methods
@@
Figure 6: Comparison of prompt produced by the state-of-the-art ORPO [31] and by UniPrompt on the GSM8K dataset.
A.13 Effect of length on performance of prompt

### v2

@@ 5.4 Results on a Real-world Task
The task of inferring if two search queries share identical intent or not arises in search and recommendation pipelines. It is challenging because it requires domain knowledge (e.g., brands and product categories), and depends on cultural and geographical biases (e.g., “cricket” in UK vs. “cricket game” in the US). So, examples are crucial for engineering a prompt that generalizes well.
We sample 200 train, 50 validation, and a separate 2527 test user queries from a real proprietary application. More details on the dataset and the one-line initial prompt are provided in Appendix A.5.
The prompt obtained using UniPrompt improves over the best manual prompt by 5.77% on the negative (rare) class, by 0.23% on the positive class, and by 1.86% overall on the test set. The learnt prompt captures the following task facets: (1) recognizing variations in names and abbreviations, and how they do not change the context; (2) recognizing brand specificity, and how even minor variations do change the context; and (3) recognizing the specificity of terms in queries, and how lack of specific terms can indicate departure of intent.
5.5 Ablations
Impact of Clustering, Inclusion of History, and Greedy Update: The results are shown in Table 6 (in Appendix A.8). We see that clustering as well as edit history components (Section 3.1) are critical for performance of UniPrompt in all the datasets. We see a major drop of $14.8\%$ in accuracy in the Ethos dataset when clustering is removed, and a $4.3\%$ drop when history component is removed. In all the datasets except GSM8K, we find clustering is more important than history. This can attributed to limited variability of question types (all grade-8 arithmetic) in GSM8K than in others.
We also find that the greedy update rule (Section 3.1.3) proves to be superior or competitive compared to beam search in relatively easier datasets — where even less exploration produces good results, greedy proves to be a more effective update rule. On the other hand, in more complex datasets like MedQA, greedy appears to be a bad strategy. We also see that clustering examples based on feedback (“Fb Clustering”) is a better strategy than clustering based on topics, except for the Ethos dataset.
Impact of Mini-batch size and Number of Clusters:
Table 3: Impact of UniPrompt’s key hyperparameters
Hyperparameter Value Accuracy
Mini-batch Size 2 84.90%
5 90.36%
8 91.15%
Number of Clusters 2 85.81%
5 90.36%
10 87.82%
First, we vary mini-batch size in UniPrompt. The results for the ARC dataset are shown in Table 3 (in Appendix A.8). With an increase in mini-batch size, we observe an increase in accuracy. That said, it is a hyperparameter, hence there will be an optimal number for each dataset. The mini-batch size affects the feedback based on the wrong examples that are obtained in each round. Next, we vary the number of clusters in UniPrompt. We find that the parameter has a clear impact on performance. We use the default choice of 5 clusters in all our experiments, which provides concise and generalizable feedback.
Impact of initial prompt and Expert model capacity: In Table 7 (Appendix A.8), we find that one-line task description initialization for UniPrompt achieves the best accuracy on three out of four datasets. On ARC, initializing with the prompt generated by the Llama2-13B model gives significant improvement over other initializations. In Table 5 (Appendix A.8), we show UniPrompt improves prompts for more capable solver LLMs while using less capable expert LLMs.
6 Conclusions
We presented a method inspired by the human prompt engineering process to generate complex prompts from scratch that include different facets of a task. Our algorithm provides significant improvements over baseline prompt generation methods on multiple standard datasets. Just like in-context learning (Ji et al., 2024), task facet learning could also benefit from connections to submodular optimization (Krause and Golovin, 2014). We leave this as future work.
7 Limitations
We provide an analysis of the impact of model size on the amenability to prompt optimization in Appendix A.1. However, in our evaluation, we only use GPT-3.5 (and in some cases GPT-4) as the solver LLM. We want to leave extensive evaluations of using open-source LLMs as solver LLMs, and perhaps even as expert LLMs, to future work. Further, we also want to evaluate on other generative tasks, besides the code generation task we study in the paper.
Appendix A Appendix
A.1 When is directional text optimization feasible?
Figure 2: Estimating (probabilistic) Lipschitz constant of models (Definition 1) on (left) Ethos (middle) GSM8K and (right) MedQA datasets for GPT-4 and GPT-3.5 models.
Consider the class of sequential algorithms like ProTeGi (Pryzant et al., 2023) and TextGrad (Hou et al., 2023) . The objective is to improve the accuracy of a given (black-box) solver LLM $f:\mathcal{X}\rightarrow\mathbb{R}$ that takes as input a prompt $\mathbf{x}\in\mathcal{X}$ and outputs the average accuracy on a validation set $D_{v}$. Since the set of prompts is combinatorially large, we assume that all prompts can be embedded in a vector space such that distance between two prompts in the space correspond to their semantic similarity. The prompt optimization problem can be written as $\arg\max_{\mathbf{x}\in\mathcal{X}}f(\mathbf{x};D_{v})$.
Previous work has shown that LLMs can be brittle to their input: changing the prompt slightly can create a significant difference in performance (Zhuo et al., 2023). We want to understand if the optimization problem is well-conditioned. Typically, conditioning can be determined by the Hessian. However, since $f$ is black-box, we approximate it by measuring sensitivity, or more specifically, Lipschitz continuity near the optimal solution. Based on prior work on defining continuity of neural networks (Mangal et al., 2020), we use a probabilistic notion.
Definition 1 (Probabilistic Lipschitz Continuity (Mangal et al., 2020)).
Given a probability distribution over inputs $\mathcal{X}$, $r\geq 0$, and a distance measure $d$ such as $\ell_{1}$ or $\ell_{2}$ norm, a function $f:\mathcal{X}\rightarrow\mathbb{R}$ is $(L,\epsilon)$-probabilistically Lipschitz with constant $L\geq 0$, if
$\Pr_{\mathbf{x},\mathbf{x}^{\prime}\sim\mathcal{X}}[$ $\operatorname{d}(f(\mathbf{x}),f(\mathbf{x}^{\prime}))\leq L\cdot\operatorname{d}(\mathbf{x},\mathbf{x}^{\prime})$
$|\ \operatorname{d}(\mathbf{x},\mathbf{x}^{\prime})\leq r]\geq 1-\epsilon.$ (1)
Note the focus on small changes in input through the parameter $r$. Intuitively, the Lipschitz property bounds the maximum change in $f$ given a small change in input prompt. Typically, the lower bound of error for any sequential optimization algorithm over $f$ is directly proportional to the Lipschitz constant $L$ (Malherbe and Vayatis, 2017). Therefore, for faster convergence, it is desirable to have a low $L$, especially near the optimal solution.
Empirically, we estimate $L$ by sampling task-relevant prompts so that they are close to the optimal solution. Then we make small changes to the prompt such that the semantic meaning stays the same and measure the change in $f$ (See Appendix A.2 for experimental details). We show the change in $f$ per change in input for GPT4 and GPT3.5 models in Figure 2 for the Ethos, GSM8K and MedQA datasets. Assuming $\epsilon=0.05$, probabilistic Lipschitz constant $L$ for GPT4 is $<1$, whereas it is higher for GPT3.5. Thus, as the model sizes increases, the probabilistic Lipschitz constant decreases. So, larger models are more amenable to prompt optimization.
A.2 Details on estimation of Lipschitz constant $L$
TO calculate the Lipschitz constant for a given LLM and task, we take a human written promp and generate it’s paraphrases using GPT-4. We prompt GPT-4 with the following text: “You are given a sentence, you have to generate 30 paraphrases of the sentence, make sure that the core content of each paraphrase is same, you can use add, subtract or change words". These paraphrases are then evaluated on the validation set $D_{v}$. For a measure of distance between two prompts, we take the cosine similarity between the embeddings of two prompts. We use text-ada-002 for generating the text embeddings for prompts.
A.3 Computational Complexity
We now consider the compute complexity of UniPrompt in terms of the number of expert or solver LLM calls per epoch, stage-wise.
Clustering: First, we evaluate all the training examples using the current prompt. Second, for every wrongly predicted example, we obtain feedback from the expert LLM. Third, for the given set of feedbacks, we use a single call to cluster it into $l$ clusters. Each of the above steps incurs $O(N)$ queries, so the total query complexity of the clustering stage is $O(N)$. Finally, for each example, i.e., (question, answer) pair, we simply map it to the $l$ clusters (no LLM calls).
Mini-batch feedback and Batch-level aggregation: At a given epoch, we evaluate every question in the mini-batch using the current prompt and the solver LLM ($N$ queries overall). Next, we obtain one feedback over all the wrong questions in the mini-batch $m$ ($N/|m|$ queries). We use one call to aggregate these feedbacks. For prompt selection, we evaluate $4$ prompts on the batch $b$ (2 per beam), so $O(4|b|)$ queries per batch. Hence overall query complexity is $N+N/|m|+4N+1$ or $O(N)$.
With LLM throughput of 0.5 qps, a training + validation set of 300 examples, 10 clusters, and 20 epochs, it takes under 7 hours to train.
A.4 SLM Training Details
To induce the ability of structured prompt generation in a smaller language model, we curate a section-wise dataset of around 12,000 task-prompt pairs. The tasks for training dataset creation were taken from tasksource library (Sileo, 2023) that contains around five hundred classification tasks. We extract the task description from tasksource-instruct, which contains tasksource dataset recasted with instructions. For instance, the task description for BIG-bench Entailed Polarity task is, "Given a fact, answer the following question with a yes or a no". The dataset provides diverse tasks and their short description, but not the human-generated prompts for each task. To approximate human-generated prompts, we use GPT-4 as a teacher model.
A.5 Data Set Creation of Real-World Task
@@
A.6 Prompt to Llama2-13B for fine-tuning
### Instruction:
You are a prompt engineer, you have
to write a structured prompt.
For the given task description,
examples and section description,
write the contents of the section
that align with
section description.
### Task Description:
{data_point[’task_description’]}
### Section Description:
{data_point[’section’]}:
{section_descriptions\
[data_point[’section’]]}
### Response:
{data_point[’prompt’]}
A.7 Prompt Initialization
@@
A.8 Additional Results
Table 4: Performance (solved problems) of UniPrompt (GPT-4-Turbo solver) on code generation datasets, compared to GPT-4 (OpenAI et al., 2023) and newer models.
Method HumanEval MBPP
GPT-4 67.0 87.5
GPT-4-Turbo 87.1 90.9
GPT-4o 90.2 92.4
UniPrompt 93.8 92.5
Table 5: Ablation of LLM choices for UniPrompt on the Ethos dataset. ‘Init’ and ‘Final’ denote initial (i.e., task description) and final prompt accuracies.
Expert LLM Solver LLM Init Final
GPT-3.5-T GPT-3.5-T 76.8 82.4
GPT-4 GPT-3.5-T 76.8 92.3
GPT-3.5-T GPT-4 89.8 91.4
GPT-4 GPT-4 89.8 94.3
Ethos ARC MedQA GSM8K
UniPrompt -- History 88.0 84.6 55.3 80.8
UniPrompt -- Clustering 77.5 82.0 54.1 81.5
UniPrompt 92.3 86.0 57.1 82.4
UniPrompt $+$ Greedy 93.7 90.5 55.5 82.3
UniPrompt $+$ Fb Clustering 87.2 91.2 58.3 82.5
Table 7: Ablation on the initial prompt for UniPrompt (best test accuracy in bold).
Init Prompt Ethos ARC MedQA GSM8K
Expert Prompt 84.0 86.0 52.3 82.4
Llama Prompt 92.0 90.5 55.5 81.5
Task Description 92.3 86.0 57.1 82.4
MedQA PubMedQA MedMCQA MMLU MG
Ours 80.9 70.3 79.2 78.0
Ours + kNN 81.0 72.2 81.4 94.0
Ours + kNN + CoT 83.9 74.7 82.6 96.0
Ours + kNN+ CoT + Ensemble 87.0 75.6 84.5 99.0
MedPrompt 80.6 71.2 79.1 98.0
@@ Introduction:
Assume the role of a science expert and answer the given question by selecting one
of the options A, B, C or D.
     1. Understand and solve science questions by selecting the best answer
     from a given list of options.
     2. Identify the logic behind the choices provided and make an informed decision.
     3. Use contextual clues to choose the most accurate answer.
     4. Be aware of the differences between science and everyday language.
@@ 4. Look for clues: Look for clues in the question and the
A.8.1 Examples of training dataset
@@ 3. To make the task more manageable, first focus on swapping the
A.9 Prompt for identifying important facets
@@ A.12 Feedback Prompts
You are a teacher and you have to give feedback to your
students on their answers.
@@
You are given a question, it’s true answer and answer given by
student. You are also given the explanations written by your
students while solving the questions.
@@
which edit can be applied to aid the student in solving the
wrong question.
@@ 5. You can add specifics like examples, definitions etc make
Your summary cannot include more than four functions.
Make sure that the content is useful,
@@ A.13 Editing Prompt
A.14 Example of prompt evolution using our method
@@
A.15 Comparision of our method with existing methods
@@
Figure 5: Comparison of prompt produced by the state-of-the-art ORPO (Yang et al., 2023) and by UniPrompt on the GSM8K dataset.
A.16 Effect of length on performance of prompt
@@
A.17 Do diverse task facets organized as sections really help?
We want to empirically validate if all the diverse task facets that UniPrompt learns indeed contribute to the performance gains that we observe in Table 1. We consider two ablations:
Table 10: Sensitivity of UniPrompt to expert LLM prompts, on the Ethos dataset.
Expert LLM Prompt for UniPrompt Test Accuracy
Simple prompt for mini-batch feedback 83.5
Simple prompt for batch feedback 91.0
Detailed prompts (Appendix A.12) 93.7
A.18 Sensitivity to prompts used for expert LLMs in UniPrompt
The prompts used for expert LLMs in our algorithm, i.e., for clustering, feedback over batches and mini-batches, and editing, do matter for obtaining good performance. However, note that the prompts are task-agnostic and can be used as-is for new tasks. Moreover, prompts for clustering and editing are very simple and involved minimal human effort. Further, to study the reliance of UniPrompt on the quality of feedback prompts, we run an ablation study, where we replace the engineered prompts for feedback at batch and mini-batch levels with simpler prompts. The results are given in Table 10 for the Ethos dataset. We observe that the performance of UniPrompt depends heavily on the prompt used for obtaining feedback at mini-batch level; whereas simplifying prompt for feedback at the batch level has much less impact on the final accuracy.
A.19 UniPrompt training behavior
An example of evolution of prompts using our algorithm is given in Appendix 3. It starts with a simple description of task and adds important facets like differentiating between hate speech and rudeness. In contrast, ProTeGi (Pryzant et al., 2023) yields a rather terse prompt on the same dataset: “Does the following text contain language that targets a group of people based on their religion, gender, or other personal characteristics?”.
Figure 6: Training curves for MedQA (left) and ARC (middle) datasets when UniPrompt is initialized with (published) state-of-the-art prompts; (right) ablation of facets on Ethos.
The training curves in Figure 6 show that our method initially performs edits on the prompt that simultaneously increase the train as well as the validation accuracy. After about 10 or 15 iterations (each batch update is an iteration), validation accuracy decreases while train accuracy continues increasing, indicating overfitting; which we overcome using early stopping.